OpenStudy (anonymous):

which one is the same as log(3x) 5y ? a. log 5x/3y b. log 3x/5y c. log(5y)/log(3x) d. log(3x)/log(5y) 3x is the base, by the way.

jimthompson5910 (jim_thompson5910):

Are you familiar with the change of base rule?

jimthompson5910 (jim_thompson5910):

If not, the change of base rule says \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\]

OpenStudy (anonymous):

not really :/

jimthompson5910 (jim_thompson5910):

What is the base in your problem?

OpenStudy (anonymous):

3x

jimthompson5910 (jim_thompson5910):

good, so b = 3x

OpenStudy (anonymous):

so logx/log3x?

jimthompson5910 (jim_thompson5910):

The argument (the stuff inside the log) is 5y, so we replace that 'x' in the formula I wrote above with '5y' to get... \[\Large \log_{3x}(5y) = \frac{\log(5y)}{\log(3x)}\]

OpenStudy (anonymous):

ohhh. okay! so is that how it always is on these types of problems?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay (: thank you!

jimthompson5910 (jim_thompson5910):

any time you have a log and you want to write in another base, then you'll use the change of base formula

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

i have one more question :P what is the solution to log3x-2 125 = 3 ? 3x-2 is the base

jimthompson5910 (jim_thompson5910):

\[\Large \log_{3x-2}(125) = 3\] \[\Large (3x-2)^3 = 125\] \[\Large 3x-2 = \sqrt[3]{125}\] \[\Large 3x-2 = 5\] I'll let you finish

OpenStudy (anonymous):

so its... 3x = x = 3/3 which x=1 ?

jimthompson5910 (jim_thompson5910):

no, you add 2 to both sides to get \[\Large 3x-2 = 5\] \[\Large 3x-2+2 = 5+2\] \[\Large 3x = 7\]

OpenStudy (anonymous):

ADD two . shoot i subtracted . my bad . so 7/3 is the answer then !

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thank yo uso much :D

jimthompson5910 (jim_thompson5910):

sure thing